# PrivacyRestore: Privacy-Preserving Inference in Large Language Models via Privacy Removal and Restoration

### v3

@@ Appendix I Hyperparameter analysis
Appendix L Supplements Of Empirical Privacy Protection
We present empirical evidence of the privacy protection capabilities of $d_{\chi}$-privacy and PrivacyRestore by implementing various attacks on these privacy-preserving methods. Lower attack performance indicates stronger privacy protection provided by these methods.
L.1 Empirical Privacy Protection of $d_{\chi}$-privacy
As presented by Feyisetan et al. (2019, 2020), the $d_{\chi}$-privacy mechanism protects input by injecting noise into the token embeddings and replacing the original tokens with their nearest neighbors. To attack the garbled text input, we implement two types of attacks: prompt injection attack (Suo, 2024) and attribute inference attack (Li et al., 2022), both commonly used for attacking text inputs. In prompt injection attack, additional instructions are added before and after the garbled text input, prompting the model to output the original text. The attack’s performance is measured by calculating the ROUGE-L score between the generated text and the original input. Attribute inference attack performs classification on the garbled text, where the target labels are the token IDs of the original input. The attack’s performance is evaluated using the classification F1 score. The details of these attack methods are presented in Appendix M.
L.2 Empirical Privacy Protection of PrivacyRestore
As stated in Section 6.3, the attack object of PrivacyRestore is the meta vector and we implement two types of attack: the embedding inverse attack (Li et al., 2023b; Morris et al., 2023) and attribute inference attack (Li et al., 2022), both commonly used methods for attacking embeddings. Embedding inverse attack utilizes the generative model to generate the privacy spans from the meta vector. Embedding inverse attack measure the attack performance by computing the ROUGE-L between the generate output and the privacy spans. Attribute inference attack performs classification on the meta vector and the target labels are the token IDS of the privacy spans. Attribute inference attack compute the F1 score to evaluate the attack performance. The details of these attack methods are presented in Appendix M.
As shown in Figure 3(c) and 3(d), the ROUGE-L score for the embedding inverse attack remains nearly stable across different $\alpha$ values in the Pri-SLJA and Pri-DDXPlus datasets. What’s a little strange is the ROUGE-L score in the Pri-NLICE dataset shows a slight increase. The possible reason is that higher ratio indicating more privacy spans consider and resulting longer reference string when compute the ROUGE-L score. Since ROUGE-L measures the overlap between the generated output and the reference string, a longer reference string may slightly boost the score. The F1 score for the attribute inference attack remains stable across all three datasets. The stable performance in both attack scenarios provides empirical support for Theorem 5.2.
Appendix M Detailed Implementation of Attack Methods
M.1 Prompt injection Attack
$d_{\chi}$-privacy injects noise into the original user inputs and transmits the garbled inputs to the server to protect the privacy spans. We employ prompt injection attack to recover the initial question from the garbled inputs. Following Perez and Ribeiro (2022); Suo (2024), we insert additional instructions before and after the user inputs to prompt the model to output the original user input rather than the normal response. The template for the additional instructions is provided in Appendix O.4.
We set the maximum generation length for the prompt injection attack to 256 tokens. To evaluate the attack’s performance, we calculate the ROUGE-L score between the generated output and the original user input. A higher ROUGE-L score indicates greater overlap between the recovered text and the original input, signifying more successful attack results.
M.2 Attribute inference attack
Attribute inference attack attempts to steal user inputs by performing classification on the garbled inputs, where the target labels correspond to the token IDs of the original inputs. Since each input contains multiple tokens, this classification task is naturally a multi-label classification problem. Following Li et al. (2022), we utilize a multi-layer perceptron (MLP) model as the classifier. The input dimension is 4096 and the output dimension is the size of the whole vocabulary size. To evaluate the attack’s performance, we calculate the F1 score of the classification, where a higher F1 score indicates a more successful attack. The attack targets can include garbled text from $d_{\chi}$-privacy, paraphrased text, or the meta vector from PrivacyRestore. The implementation details for text and vectors may vary slightly.
Attribute inference attack on meta vector. To attack the meta vector from PrivacyRestore, we can directly use a fully-connected layer to transform the meta vector’s dimension from $128\times K$ to the classifier’s input dimension 4096. We then perform classification on the transformed meta vector.
Attribute inference attack on text. For garbled text from $d_{\chi}$-privacy or paraphrased text, we first transform the text into a vector representation. We utilize Llama2-chat-7b to process the text input and obtain the last token’s hidden state as the vector representation. Classification is then performed on this hidden state.
M.3 Embedding inverse attack
Different from attribute inference attack, embedding inverse attack steal the user inputs through the generative model to generate the original user inputs. We utilize the GPT-2 model (Radford et al., 2019) as the generative model and set the maximum generation length to 256. We finetune the GPT-2 model on the training set for 20 epoch using the learning rate of 1e-5. To evaluate the attack’s performance, we compute the ROUGE-L score between the generated output of the GPT-2 model and the original user input, where higher scores indicate better attack effectiveness. Similar to the attribute inference attack, the implementation of the embedding inverse attack differs between text and meta vectors.
Embedding inverse attack on meta vector. We use a fully-connected layer to transform the meta vector’s dimension to the dimension of hidden state of GPT-2 model. Then we directly input the transformed meta vector as the input embedding.
Embedding inverse attack on text. We use Llama-2-chat-7b to process the text input and extract the last token’s hidden state as the vector representation. This vector is then transformed by a fully connected layer to match the hidden state dimension of the GPT-2 model. Finally, it is fed into the GPT-2 model as the input embedding for subsequent generation.
Appendix N Example Outputs of PrivacyRestore
@@
Appendix O Prompt Template Details
O.1 Classification of Privacy Spans.
O.1.1 Medical Datasets (Pri-DDXPlus/Pri-NLICE).
Prompt template shown in Figure 7 is for GPT and used to classify symptoms in Pri-DDXPlus/Pri-NLICE dataset into sensitive and non-sensitive categories. GPT grades the symptoms on a scale of one to five based on sensitivity, with levels greater than three considered private spans in the Pri-DDXPlus/Pri-NLICE dataset.
O.1.2 Legal Dataset (Pri-SLJA).
Prompt template shown in Figure 8 is for GPT and used to classify the case details in Pri-SLJA dataset into sensitive and non-sensitive categories. GPT grades the symptoms on a scale of one to five based on sensitivity, with levels greater than three considered private spans in the Pri-SLJA dataset.
O.2 Generation Prompts.
O.2.1 Medical Datasets (Pri-DDXPlus/Pri-NLICE).
Prompt template shown in Figure 9 is for Llama model and used during model generation for the Pri-DDXPlus/Pri-NLICE datasets.
O.2.2 Legal Dataset (Pri-SLJA).
Prompt template shown in Figure 10 is for Llama model and used during model generation for the Pri-SLJA datasets.
O.3 Evaluation of Generated Output.
O.3.1 Medical Datasets (Pri-DDXPlus/Pri-NLICE).
@@
O.3.2 Legal Dataset (Pri-SLJA).
@@
O.4 Prompt Injection Attack.
Prompt template shown in Figure 13 is for Llama model and used to carry out a prompt injection attack, translating the garbled text back into the original text.

### v5

@@ Appendix X Privacy Spans Over-Removal
Appendix Y Prompt Template Details
Y.1 Classification of Privacy Spans.
Y.1.1 Medical Datasets (Pri-DDXPlus/Pri-NLICE).
Prompt template shown in Figure 9 is for GPT and is used to classify symptoms in Pri-DDXPlus/Pri-NLICE dataset into sensitive and non-sensitive categories. GPT grades the symptoms on a scale of one to five based on sensitivity, with levels greater than three considered private spans in the Pri-DDXPlus/Pri-NLICE dataset.
Y.1.2 Legal Dataset (Pri-SLJA).
Prompt template shown in Figure 10 is for GPT and is used to classify the case details in Pri-SLJA dataset into sensitive and non-sensitive categories. GPT grades the symptoms on a scale of one to five based on sensitivity, with levels greater than three considered private spans in the Pri-SLJA dataset.
Y.2 Rewriting of User Queries.
Y.2.1 Medical Datasets (Pri-DDXPlus/Pri-NLICE).
The prompt template shown in Figure 11 is designed for GPT and is utilized to rewrite medical queries in the Pri-DDXPlus and Pri-NLICE datasets.
Y.2.2 Legal Dataset (Pri-SLJA).
The prompt template shown in Figure 12 is designed for GPT and is utilized to rewrite medical queries in the Pri-SLJA dataset.
Y.3 Generation Prompts.
Y.3.1 Medical Datasets (Pri-DDXPlus/Pri-NLICE).
Prompt template shown in Figure 13 is for Llama model and is used during model generation for the Pri-DDXPlus/Pri-NLICE datasets.
Y.3.2 Legal Dataset (Pri-SLJA).
Prompt template shown in Figure 14 is for Llama model and is used during model generation for the Pri-SLJA datasets.
Y.4 Evaluation of Generated Output.
Y.4.1 Medical Datasets (Pri-DDXPlus/Pri-NLICE).
@@
Y.4.2 Legal Dataset (Pri-SLJA).
@@
Y.5 Prompt Injection Attack.
Prompt template shown in Figure 17 is for Llama model and is used to carry out a prompt injection attack, translating the garbled text back into the original text.
@@
Figure 11: Prompt template for rewriting queries in medical datasets
Figure 12: Prompt template for rewriting queries in legal datasets